OpenStudy (anonymous):

A dehydrated patient needs a 6.1% saline IV. Unfortunately, the hospital only has bags of 5% and 7% saline solutions. How many liters of each of these solutions should be mixed together to yield 2 liters of the desired concentration?

OpenStudy (dumbcow):

not quite..... x+y = 2, 5x+7y = (6.1)(2)

OpenStudy (anonymous):

now i need to solve it

OpenStudy (anonymous):

thats where i get confused

OpenStudy (dumbcow):

yes use substitution...........y=2-x............plug it in for y.............5x+7(2-x) = 12.2

OpenStudy (anonymous):

so thats the answer

OpenStudy (dumbcow):

whats the answer? .................the goal is to find the number of liters that must be mixed together, you need to have x = ? and y = ?, before you're done

OpenStudy (anonymous):

x=5and y =7

OpenStudy (dumbcow):

does 5+7 = 2 ?

OpenStudy (anonymous):

no

OpenStudy (dumbcow):

sorry i gotta go............work on solving this equation for x: 5x+7(2-x) = 12.2

OpenStudy (anonymous):

thanks